OpenStudy (romantic_ch3micals):

Help. Medal Graph this function by translating y = IxI y = IxI -6

OpenStudy (danjs):

shift the whole thing down by 6 units

OpenStudy (romantic_ch3micals):

i dont get it

OpenStudy (romantic_ch3micals):

@DanJS

OpenStudy (danjs):

any value you put in for x, if you change the graph to \[y =\left| x \right| - 6\] the y value will be 6 less

OpenStudy (danjs):

say x = 5 Normally y = 5, but with this new -6 term, you get y=5-6 = -1 it is shifted down 6 units

OpenStudy (danjs):

for any value of x you choose

OpenStudy (danjs):



OpenStudy (romantic_ch3micals):

thanks

OpenStudy (danjs):

welcome